### v1

# EVA-Score: Evaluation of Long-form Summarization on Informativeness through Extraction and Validation

Abstract
Summarization is a fundamental task in natural language processing (NLP) and since large language models (LLMs), such as GPT-4 and Claude, come out, increasing attention has been paid to long-form summarization whose input sequences are much longer, indicating more information contained. The current evaluation metrics either use similarity-based metrics like ROUGE and BERTScore which rely on similarity and fail to consider informativeness or LLM-based metrics, lacking quantitative analysis of information richness and are rather subjective. In this paper, we propose a new evaluation metric called EVA-Score using Atomic Fact Chain Generation and Document-level Relation Extraction together to automatically calculate the informativeness and give a definite number as an information score. Experiment results show that our metric shows a state-of-the-art correlation with humans. We also re-evaluate the performance of LLMs on long-form summarization comprehensively from the information aspect, forecasting future ways to use LLMs for long-form summarization.
EVA-Score: Evaluation of Long-form Summarization on Informativeness through Extraction and Validation
Yuchen Fan1thanks:  The work was done when he worked as an intern at Tsinghua University, China., Xin Zhong1, Chengsi Wang1, Gaochen Wu2, Bowen Zhou2thanks:  Corresponding authors 1 Beijing University of Posts and Telecommunications 2 Department of Electronic Engineering, Tsinghua University yuchenfan48@gmail.com zhoubowen@tsinghua.edu.cn
1 Introduction
Figure 1: Examples show the dilemma of ROUGE. The blue annotation is the atomic information contained in the text. Even though Candidate Summarization 1 is rewarded with a higher ROUGE Score, it contains less information than Candidate Summarization 2. It inspires us that ROUGE and BERTScore are not reasonable for evaluating summarization.
Summarization has long been a fundamental task in NLP and since LLMs came out (OpenAI et al., 2024; Touvron et al., 2023), many research leverages LLMs for summarization (Zhang et al., 2023b; Wang et al., 2023b). More recently, works on summarization moved to Long-form scenarios where the input sequence length is extremely larger, resulting in more information contained in it. Traditional evaluation metrics may be not suitable for evaluation of long-form summarization due to the great informativeness, thus calling for new evaluation metrics.
Past evaluation metrics for summarization include some static evaluation methods, like ROUGE-k and ROUGE-l (Lin, 2004) which calculate the k-gram overlap or the longest common substring between the reference and the candidate. With the development of pre-trained language models, BERTScore (Zhang et al., 2020) is widely used as a new evaluation metric based on pre-trained word embeddings. More recently, some works have begun to pay attention to evaluating long-form summarization (Kim et al., 2024; Chang et al., 2024) which uses prompts to define criteria to assign different scores for different aspects, like fluency, consistency, and coherence. However, traditional and LLM-based evaluation metrics face their dilemmas respectively.
Figure 2: Pipeline of our evaluation metric. We begin with atomic fact generation. For generated facts, we use LLMs to check whether they can be treated as a logic chain and use Document-level Relation Extraction as a complement. Finally, when analyzing the truthfulness of a given fact, we use retrieval to find several reference facts and provide previous states.
For Similarity-based Metrics Since they are based on similarity, they cannot represent informativeness and are unsuitable in such an LLM era, where responses from LLMs are flexible and contain many synonyms and different syntax. For LLM-based Metrics They are insensitive to nuanced differences existing in summaries and their quantitative results are not reliable even if given detailed criteria for evaluation due to the black-box property.
We propose a new evaluation metric whose result is an objective and quantitative analysis but can fully catch the information contained in the summary. We call this metric Extraction-and-Validation Score. We begin with extracting atomic facts from reference and candidate summaries. Different from FActScore (Min et al., 2023), we claim that some information cannot live without a context, so we treat them as part of logic chains. Afterward, we introduce Document-level Relation Extraction to remedy the loss of hierarchical relations. Then for each atomic fact to be verified, we use information retrieval to find the most relevant atomic facts as references and prompt LLMs to focus only on newly added information and decide whether it is entailed or not. Intuitively, LLMs for evaluation can be more accurate in single verification than in complex situations. We ensure that only one piece of information is evaluated at a time, simplifying the validation process. And by introducing Document-level Relations, we form a more comprehensive information set, maintaining the completeness of the evaluation.
We conduct comprehensive experiments by comparing EVA-Score with current evaluation metrics, including traditional and LLM-based. The results show that EVA-Score achieves the best human alignment compared to SOTA past evaluation metric, BERTScore by 0.166, 0.120, and 0.09 in Pearson, Spearman, and Kendall correlation respectively at the text level. At the system level, it shows human-comparable evaluation, with 0.975 in Pearson correlation and 1.0 in both Spearman and Kendall correlation. We also re-evaluate long-form summarization based on EVA-Score and the experiment results show the informativeness of different LLMs on long-form summarization. The results show the great performance of GPT-4 when the context becomes longer, ranking first in arXiv and Gov Report and second in Pubmed and BookSum. It also shows the promising performance of Mixtral-7B-Instruct for long-form summarization.
Our main contributions are as follows:
  • We propose a new evaluation metric that can generate high-quality evaluation results of a given summary, based on atomic fact chain generation and document-level relation extraction, by providing a numerical score and steps to obtain this score.
  • We conduct comprehensive experiments on several long-form summarization datasets, with the number of tokens of each record ranging from 1k to 14k. Experiment results show that our evaluation metric shows superiority in every correlation with humans.
  • We re-evaluate long-form summarization concerning informativeness which is the first to evaluate from the information aspect as far as we know. In our experiments, the state-of-the-art model GPT-4 still performs better than any other models in long context settings but lags behind some assistant models in short context settings.
2 Related Work
2.1 Long-form Summarization
Long-form summarization requires language models to understand the purpose of the context and capture core information from a long context which is a struggle for language models. During the last few years, much work from different fields regarding long-form summarization has been studied such as news reports (See et al., 2017), government reports (Cohan et al., 2018), academic papers and books (Kryściński et al., 2022), posing great challenges for language models (Lewis et al., 2019; Raffel et al., 2023). Since ChatGPT and GPT-4 (OpenAI et al., 2024) come out, LLMs show great potential to deal with long-form summarization (Lu et al., 2023; Zhang et al., 2023a; Wang et al., 2023a) because of their huge amount of pre-training corpus. However, when the context length becomes much larger, LLMs still have trouble because of limitations for input sequences and knowledge forgetting when reading a long context. More recently, some research aims at exploring LLMs which can handle longer context paving the way for long-form summarization. (Su et al., 2023; Li* et al., 2023) Yet, summarizing report-level or book-level texts is still a difficult task. (Chang et al., 2024)
2.2 Evaluation of Summarization
Tradition evaluation metrics for summarization include ROUGE variants, like ROUGE-k and ROUGE-L and BERTScore (Zhang et al., 2020). However, they only reflect the similarity between candidates and references without the information contained in the prediction which is not enough for us to choose a better summary or quantize the informativeness. Currently, several evaluation metrics (Zheng et al., 2024; Ke et al., 2023) based on LLMs have been proposed devoted to aligning model evaluation with human evaluation. Some works apply LLMs to the evaluation of long-form summarization, for example, Luo et al. (2023) uses ChatGPT to evaluate factual inconsistency for text summarization. More Recently, (Jung et al., 2024) calculates the saliency of a summary using MSM(Masked Language Model) to find how a candidate summary can be recovered given a reference summary. However, we claim that making use of the inherent ability of LLMs only is not enough to measure the ability of models to summarize for the lack of evaluation of the degree of information contained. Moreover, leveraging pre-trained probability distribution is not good enough because of its unexplainability. We hope to find a reasonable and explainable method for evaluation.
3 Preliminaries
In this section, we provide definitions for elements used in our evaluation pipeline.
Our work begins with extracting atomic facts from documents. Specifically, we denote reference summary as $s$ and candidate summary as $\hat{s}$. The extracted atomic fact lists are $S$ and $\hat{S}$ with $s_{i}$ and $\hat{s_{i}}$ being single atomic facts in the list. After atomic fact chain generation, we obtain lists containing logic chains $C$ and $\hat{C}$ for reference and candidate, while the elements of them are defined as $C_{i}$ and $\hat{C_{i}}$ and atomic fact in logic chains are defined as $C_{i_{j}}$ and ${\hat{C_{i_{j}}}}$. Finally, we conduct Document-level Relation Extraction(DocRE) to complement the original atomic fact list with hierarchical information. The final lists are $DocS$ and $Doc\hat{S}$ separately.
4 Methods
In this section, we introduce our proposed pipeline to evaluate informativeness in long-form summarization by extraction and validation. Given $s$ and $\hat{s}$, our automatic evaluation metric operates by 1) obtaining all atomic facts from $s$ and $\hat{s}$; 2) reforming atomic facts to form logic chains; 3) extracting all document-level relations and filter the duplicated one; 4) using LLM to validate whether it is true or false.
Figure 3: Examples show the generation of atomic facts. It uses demonstration to help LLMs understand the task.
4.1 Atomic Fact Generation
Atomic facts are defined as concise sentences that convey a single piece of information. In any given sentence within a model prediction, there may be multiple pieces of information that require individual verification. Consequently, evaluating the truthfulness of sentences on a one-by-one basis renders the true-or-false outcomes inadequate for assessing informativeness. Previous studies find that splitting candidate prediction into atomic facts (Min et al., 2023) improves the evaluation by a large margin and enables a more fine-grained evaluation. Following their work, we use ChatGPT with same prompt to generate atomic facts from both $S$ and $\hat{S}$ as $\left[s_{1},s_{2},s_{3}...\right]$ and $\left[\hat{s_{1}},\hat{s_{2}},\hat{s_{3}}...\right]$ where $s_{i}$ is atomic fact from reference summarization and $\hat{s_{i}}$ is atomic fact from candidate summarization.
4.2 Atomic Fact Chain Generation
Although claimed atomic, an examination of the “atomic facts” presented reveals that they often encompass more than one atomic fact. As illustrated in Figure 3, certain information cannot be independently extracted due to its reliance on contextual semantics. Consequently, in cases of cascaded atomic facts, where subsequent facts incorporate or entirely encompass information from preceding facts, if any of the preceding facts are incorrect, all subsequent facts are deemed incorrect, even if the newly added information is accurate.
To clarify, consider the examples: He played for Berekum Arsenal” and He played for Berekum Arsenal before moving to the United States.” These are two cascaded facts, with the latter not being atomic. However, the latter sentence cannot be decomposed into atomic facts because the adverbial clause depends on the entirety of the sentence for its meaning. Consequently, if the initial statement is incorrect, any subsequent statements, including those with additional correct information, are also considered incorrect.
We propose using LLMs, e.g. Mixtral-7B-instruct in our work to find logical relationships between atomic facts and treating atomic facts as a list of logic chains. The reformed lists are represented as $[C_{1},C_{2},C_{3}...]$ and $[\hat{C_{1}},\hat{C_{2}},\hat{C_{3}}...]$.
4.3 Document-level Relation Extraction
Extracting atomic facts from text primarily captures sentence-level relations. However, it is crucial to also consider document-level relations, which are challenging to discern solely through atomic fact generation. These relations are integral to understanding the contextual logic of a document. We complement Document-level relation extractions(DocRE) using the state-of-the-art model GPT-4 with in-context learning for both $s$ and $\hat{s}$ and the results are $[(h_{1},r_{1},t_{1}),(h_{2},r_{2},t_{2})...]$ and $[(\hat{h_{1}},\hat{r_{1}},\hat{t_{1}}),(\hat{h_{2}},\hat{t_{2}},\hat{r_{2}})...]$ where $(h_{i},r_{i},t_{i})$ and $(\hat{h_{i}},\hat{r_{i}},\hat{t_{i}})$ are head entity, relation and tail entity correspondingly. We then paraphrase the triples to make them natural language for post-processing.
To ensure that no sentence-level relations are included and contaminate work later, we use BERTScore along with TF-IDF to filter out the similar ones existing both in the atomic fact list and document-level relation list.
4.4 LLM for Validation
For each atomic fact $s_{i}$ in the candidate atomic fact list, we retrieve $k$ most relevant atomic facts from the reference summarization atomic fact list where k is a hyper-parameter which is a trade-off between accuracy and efficiency. For each atomic fact $\hat{C_{i_{j}}}$ in the corresponding logic chain $\hat{C_{i}}$, we check out one by one according to $s_{i}$ from $k$ most relevant atomic facts in $DocS$ represented as $f(\hat{C_{i_{j}}},s_{i})$. Specifically, we use Mixtral-7B-Instruct as our validation model. By providing the validation model with previous states and prompting it not to pay attention to previous information, we ensure that one piece of information is treated at a time.
We calculate our Precision and Recall as follows:
$\text{Precision}=\frac{1}{|Doc\hat{S}|}\sum_{i=0}^{n}\mathds{1}\left(f(\hat{C_{i_{j}}},s_{i})\right)$ (1)
$\text{Recall}=\frac{1}{|DocS|}\sum_{i=0}^{n}\mathds{1}\left(f^{-1}(s_{i},\hat{C_{i_{j}}})\right)$ (2)
where $n$ is the number of atomic facts in $Doc\hat{S}$. We define EVA-Score as the harmonic mean of Precision and Recall.
5 Experiment
5.1 Dataset
To test the applicability of EVA-Score, we collect long-form summarization with different context lengths. Specifically, we adopt CNN Dailymail(See et al., 2017), PubMed, GovReport(Cohan et al., 2018), arXiv, BookSum(Kryściński et al., 2022). In particular, we use the chapter-level summarization version of BookSum instead of the book level for less difficulty. For all these datasets, we use validation datasets for evaluation.
Dataset CNN Dailymail Pubmed arXiv Gov Report BookSum
Average Text Length 781 3111 5894 7797 14337
Average Summary Length 56 216 172 573 570
Table 1: The Average Text Length and the Average Summary Length in our experimented dataset.
5.2 Model
To perform a fair comparison, we experiment on a wide range of Generative Language Models of different numbers of parameters, and from different architectures. For open-source models, we test EVA-Score on Mistral-7B(Jiang et al., 2023) which is a MoE model, variants of Llama-2-chat(Touvron et al., 2023), with parameters 7B and 13B, variants of Vicuna(Zheng et al., 2023) with parameters 7B and 13B and Gemma-7B-it(Team et al., 2024). For close-source models, we experiment on ChatGPT and GPT-4(0125-preview version).
5.3 Baseline
We compare the results of EVA-Score with traditional evaluation metrics like ROUGE and BERTScore(Zhang et al., 2020) which are based on text overlap and word embeddings. Recently, many works have leveraged LLMs to obtain the informativeness of a given text corresponding to a reference. We include Vicuna-13B and Mixtral-7B-Instruct as our open-source LLM evaluators and ChatGPT and GPT-4(0125-preview version) as our close-source LLM evaluators. For all the LLMs, we instruct them using the prompt from Table 12, imitating the human annotation process.
5.4 Human Annotation
We are committed to developing an evaluation metric that resonates with human judgment. To achieve this, we randomly select 30 records from the dataset and assess the informativeness of each record by ourselves. Initially, we identify sentence-level and document-level atomic facts, denoted as $S$ and $\hat{S}$, from $s$ and $\hat{s}$, respectively. They then determine the overlap between $S$ and $\hat{S}$. Redundant information, which appears multiple times, is recorded only once. Subsequently, we calculate Precision and Recall for these overlaps and derive the F1 Score as their harmonic mean. Furthermore, we ensure that each record is evaluated by at least two annotators to maintain assessment fairness.
5.5 Main Results
Level Text-Level System-Level
Metric $\sigma$ $\gamma$ $\tau$ $\sigma$ $\gamma$ $\tau$
Traditional Evaluation Metric
ROUGE-1 0.305 0.160 0.114 0.455 0.800 0.667
ROUGE-2 0.413 0.203 0.127 0.014 0.399 0.333
BERT-Score 0.534 0.533 0.413 0.623 0.600 0.333
Model-based Evaluation Metric
Mixtral-7B-Instruct 0.097 0.052 0.030 -0.653 -0.400 -0.333
Vicuna-13B 0.065 -0.040 -0.029 0.367 0 0
ChatGPT 0.138 0.210 0.162 0.430 0.600 0.333
GPT-4 0.308 0.329 0.229 0.505 0.600 0.333
EVA-Score (Ours) 0.710 0.673 0.503 0.975 1.000 1.000
Table 2: Text-level and system-level Pearson($\sigma$), Spearman($\gamma$), and Kendall($\tau$) correlations with different evaluation metrics. The highest correlation among the methods is bold
In this section, we report the correlation between each evaluation metric and human annotations. We adopt text-level and system-level Pearson($\sigma$), Spearman($\gamma$), and Kendall($\tau$) correlation. Text-level correlation is calculated by computing the correlation coefficients between human evaluations and automatic evaluation metrics and then averaging these coefficients for all generated texts per instruction. System-level correlation, on the other hand, is determined by computing the correlation coefficients between human evaluations and automatic evaluation metrics, but this time for the average score of each dataset.
From Table 2, we observe that our EVA-Score outperforms all other model-based and traditional evaluation metrics in achieving the highest correlation with human judgments across all indicators at both levels. At the System Level, EVA-Score achieves perfect Spearman and Kendall correlations of 1.00, demonstrating complete alignment with human assessments. In contrast, Traditional Evaluation Methods, while somewhat correlated with human evaluations, generally lack strong correlations, highlighting their limitations.
Surprisingly, we find that the information scores assigned by Vicuna-13B show little correlation with humans, and At the System Level, the scores from Mistral-7B show even negative correlation with humans. We attribute this to the situation that the informativeness is obtained by quantitative calculation using EVA-Score and Human Annotation. The instruction may be too complex for them to deal with within a single interaction which requires them first to extract atomic facts and document-level relations and verify all of them one by one. However, for models like ChatGPT and GPT-4, while their correlation with human judgments at the Text Level remains minimal, they exhibit promising outcomes at the System Level using only In-Context Learning.
We also report our results on several long-form summarization datasets, including CNN Dailymail, PubMed, GovReport, arXiv, and BookSum, using three different evaluation metrics, i.e. EVA-Score, ROUGE-1, and BERTScore. Since this work focuses on the evaluation of long-form summarization, we don’t pursue feeding long context into LLMs without information loss. Instead, we believe that useful information mostly appears at the beginning or the end of a text, we just simply truncate the context from both the beginning and end to meet the models’ max input sequence length.
Dataset CNN Dailymail Pubmed arXiv Gov Report BookSum
Metric EVA R-1 BS EVA R-1 BS EVA R-1 BS EVA R-1 BS EVA R-1 BS
Vicuna-7B 0.501 0.441 0.873 0.327 0.398 0.832 0.327 0.344 0.838 0.239 0.308 0.846 0.145 0.212 0.825
Vicuna-13B 0.462 0.432 0.872 0.431 0.388 0.837 0.346 0.348 0.832 0.154 0.298 0.849 0.144 0.188 0.827
Mixtral-7B-Instruct 0.499 0.422 0.872 0.447 0.447 0.847 0.444 0.448 0.845 0.313 0.399 0.854 0.259 0.377 0.839
Llama-2-7B-chat 0.399 0.386 0.868 0.358 0.421 0.845 0.411 0.435 0.851 0.193 0.198 0.798 0.186 0.268 0.826
Llama-2-13B-chat 0.453 0.423 0.868 0.416 0.405 0.841 0.444 0.332 0.830p 0.296 0.378 0.857 0.154 0.280 0.827
Gemma-7B-it 0.431 0.383 0.858 0.492 0.417 0.842 0.385 0.359 0.831 0.250 0.291 0.843 0.154 0.256 0.818
ChatGPT 0.472 0.373 0.864 0.444 0.359 0.838 0.443 0.368 0.838 0.302 0.261 0.854 0.182 0.202 0.826
GPT-4 0.466 0.376 0.863 0.454 0.407 0.838 0.453 0.361 0.835 0.325 0.332 0.854 0.257 0.261 0.824
Table 3: The results of Several LLMs on different long-form summarization datasets. We denote EVA-Score as EVA, R-1 as ROUGE-1, BS as BERTScore. All the values are normalized between 0 and 1. We indicate the highest score of EVA-Score using bold.
Referring to Table 3, we observe that as the context lengthens, all tested models experience varying degrees of performance decline. Particularly notable is the performance at the part-of-book level, i.e. chapter summarization in BookSum, where the results are markedly poor compared to the ground truth. This issue, however, can be mitigated by employing methods that feed the entire context into the model, rather than truncating it as we have done in our initial experiments. Additionally, compared to other evaluation metrics, the EVA-Score demonstrates a more pronounced and consistent decline in performance as context length increases, with fewer fluctuations. This pattern aligns well with human cognitive processes, indicating that as information load increases, the challenge of maintaining high summarization quality also grows. This consistent pattern of decline not only reflects realistic expectations of model performance but also underscores the effectiveness of the EVA-Score in capturing these dynamics, further validating its utility from another perspective.
Moreover, models using ROPE(Su et al., 2023) or other methods, enabling longer input sequences show greater performance in long-form summarization when the context length becomes larger, mainly due to their ability to comprehend a long context or search in a long context. However, when the context is not so long, they show comparable or inferior performance with assistant models with normal sequence length, e.g. 4096 or 8192.
Under the evaluation framework of the EVA-Score, GPT-4 demonstrates robust capabilities, particularly with longer input sequences, compared to other models. In the CNN Dailymail dataset, it trails behind models such as Vicuna-7B and ChatGPT. This is primarily due to GPT-4’s tendency to generate more extensive summaries than its counterparts in settings with relatively short contexts, thereby increasing its count of atomic facts. While this might allow it to cover more points than other models, it often results in lower precision due to the inclusion of information that extends beyond the reference summary. Nevertheless, in other datasets such as arXiv and Gov Report, GPT-4 outperforms all competitors. It also secures a second-place ranking in Pubmed and BookSum, affirming its effectiveness across diverse text summarization tasks.
To be mentioned, Mixtral-7B-Instruct, compared with other models with the same parameter or even larger, shows great potential to fulfill long-form summarization and even beats GPT-4 by a small margin in chapter-level BookSum, which meets the findings of Chang et al. (2024).
5.6 Analysis of Summarization Length
Dataset CNN Dailymail Pubmed arXiv Gov Report Booksum
Indicator Avg Length Avg Length Avg Length Avg Length Avg Length
Vicuna-7B 108 214 133 258 124
Vicuna-13B 106 146 205 188 99
Mistral-7B-Instruct 109 207 212 353 251
Llama-2-7B-chat-hf 118 143 108 140 185
Llama-2-13B-chat-hf 124 184 221 236 179
Gemma-7B-it 98 139 140 177 151
ChatGPT 102 118 124 137 107
GPT-4 153 227 272 242 212
Table 4: The average summarization length(denoted by Avg Length by different model. The longest is marked underline
To elucidate why GPT-4 underperforms on the EVA-Score with a shorter context, we conduct an analysis focused on the summarization lengths produced by different models, as detailed in Table 4. Models pre-trained with longer input sequences tend to generate more extensive responses. Specifically, GPT-4 produces summaries of 153 tokens in CNN Dailymail, 227 tokens in Pubmed, and 272 tokens in arXiv and Mixtral-7B-Instruct outputs 353 tokens in Gov Report and 251 tokens in BookSum whose maximum input sequences length are 128000 and 32768 respectively.
Upon manual examination, we observe that shorter contexts typically result in shorter ground truth. Consequently, if a model generates a lengthy prediction, this verbosity is penalized with a reduced EVA-Score due to diminished precision. As the context lengthens, references become richer in information. However, these models tend to maintain a similar response length, lacking significant information addition, thereby failing to capture enough relevant points, especially in longer predictions. This discrepancy explains GPT-4’s shortfall in CNN Dailymail and the superior performance of Mistral-7B-Instruct, despite having a comparable number of parameters.
Furthermore, GPT-4, which is refined through Reinforcement Learning from Human Feedback (RLHF), typically caps its responses to summarizations under 300 tokens, even in contexts of substantial length. This behavior is a contributing factor to GPT-4’s underachievement in BookSum, where more extensive and detailed responses are crucial for capturing the full scope of the content.
5.7 Ablation Study
We report the effect of Atomic Fact Chain Generation and Document-level Relation Extraction on the number of atomic facts to be verified.
Settings AFG only AFCG AFCG + DocRE
CNN Dailymail 25 19 24
Pubmed 38 25 30
arXiv 33 22 26
Gov Report 44 26 30
BookSum 36 27 31
Table 5: Number of facts(or fact chains) to be evaluated under three settings, averaging 9 models we used for experiments, i.e. Atomic Fact Generation(AFG), Atomic Fact Chain Generation(AFCG) and Atomic Fact Chain Generation Complemented with DocRE(AFCG + DocRE).
From Table 5, datasets with longer context and richer information tend to result in more facts, and by using Atomic Fact Chain Generation, we reduce the number of cascaded facts a lot. By doing so, minor mistakes in one piece of information will not affect others. Also, DocRE complements AFG with document-level relations, which is helpful for the evaluation of the hierarchical information in the summary.
5.8 Meta Evaluation of Human Annotation
By observing the results of human annotations, we find that ground truth prefers to summarize more elegantly, using core arguments from the original context in addition to examples supported with other a little bit trivial arguments. However, LLMs try to list all of the arguments, ignoring the examples or data in the context, leading to differences with ground truth. However, there does exist some verbosity in the ground truth, indicating the natural drawback of the summarization dataset.
5.9 Analysis of Different LLM Verifier
Since our LLM Verifier should consider previous states in a logic chain and verify whether the current atomic fact is true or not, it is a rather hard problem compared with simply verifying. Intuitively, with more capable models, we will get a more human-aligned evaluation metric. We report the results using ChatGPT for a simple proof.
Metric $\sigma$ $\gamma$ $\tau$
EVA-Score-Mistral-7B-Instruct 0.710 0.673 0.503
EVA-Score-ChatGPT 0.737 0.672 0.543
Table 6: Human correlation using ChatGPT as EVA-Score LLM verifier compared with EVA-Score with Mistral-7B-Instruct. The larger one is denoted is bold.
Leveraging ChatGPT results in more human correlation in Pearson and Kendall correlation and comparable Spearman correlation, which is consistent with our recognition. However, using ChatGPT results in much more costs than open-source LLMs because of the great number of validations.
5.10 Case Study
To provide a more intuitive and straightforward insight into the pipeline of EVA-Score, we provide a detailed case in Appendix B. From the case, we can see that EVA-Score provides not only a quantitative and objective information score but also provides a human-like explanation.
6 Conclusion and Future Work
In this work, we propose a new evaluation metric called EVA-Score to evaluate informativeness in long-form summarization. It extracts atomic facts which are then reformed as logic chains and document-level relations from reference and candidate and uses LLMs for validation one by one. Experiment results show that compared with other traditional metrics or LLM-based metrics, our work shows the best human correlation as well as provides an objective explanation. Moreover, we re-evaluate long-form summarization based on EVA-Score, indicating a great challenge in book-level summarization.
We believe that EVA-Score is meant to calculate information overlap between reference and candidate, thus it can be applied to broader scenarios rather than long-form summarization only. In future work, we suggest: 1) Leveraging EVA-Score to evaluate more reference-based and information-rich tasks and 2) Utilizing EVA-Score for the generation of more informative texts.
Limitations
EVA-Score can be time-consuming to run
EVA-Score includes several pre-processing methods. We should first extract atomic facts and document-level relations two summaries and use LLMs for paraphrasing to make the relations more natural. The pre-processing takes around two minutes per record. Then for each atomic fact in the candidate, we select $k$ most relevant from the reference and use LLMs for validation. Since the number of atomic facts takes quite a lot of time. We report the time consumed correlated to the sum of the length of the candidate and reference.
Length Interval/token 0-300 300-600 600-900 900-1200
Time Consuming/s 195 402 578 629
Table 7: Time consumed regarding to different summary length
From Table 7, for a summary of relatively short length, it takes about three minutes to finish the evaluation, while for a longer summary, it may take about ten minutes which is a long time compared to traditional methods.
EVA-Score only considers Informativeness
Some existing evaluation metrics consider many other factors, like faithfulness, and consistency. EVA-Score may be combined with other evaluation metrics to provide a more comprehensive evaluation of long-form summarization.
EVA-Score Leverages GPT-4 for Doc-RE
EVA-Score implements Document-level Relation Extraction for the lack of hierarchical information. However, Doc-RE by GPT-4 maybe incomplete because GPT-4 may fail to capture the core meaning of the document level due to the difficulty of such a task. Leveraging more capable models that are fine-tuned on Doc-RE may be more helpful.
Acknowledgments
This work is supported by the National Science and Technology Major Project(2023ZD0121403). We extend our gratitude to the anonymous reviewers for their insightful feedback.
Appendix A Instructions for EVA-Score
A.1 Instruction for Atomic Fact Generation
We use the following prompt to generate atomic facts.
Instruction: Please breakdown the following sentence into independent facts:
Domonstrations
Example 1
He made his acting debut in the film The Moon is the Sun’s Dream (1992), and continued to appear in small and supporting roles throughout the 1990s.
- He made his acting debut in the film.
- He made his acting debut in The Moon is the Sun’s Dream.
- The Moon is the Sun’s Dream is a film.
- The Moon is the Sun’s Dream was released in 1992.
- After his acting debut, he appeared in small and supporting roles.
- After his acting debut, he appeared in small and supporting roles throughout the 1990s.
Example 2 In 1963, Collins became one of the third group of astronauts selected by NASA and he served as the back-up Command Module Pilot for the Gemini 7 mission.
- Collins became an astronaut.
- Collins became one of the third group of astronauts.
- Collins became one of the third group of astronauts selected.
- Collins became one of the third group of astronauts selected by NASA.
- Collins became one of the third group of astronauts selected by NASA in 1963.
- He served as the Command Module Pilot.
- He served as the back-up Command Module Pilot.
- He served as the Command Module Pilot for the Gemini 7 mission.
Table 8: The instruction for atomic fact generation
A.2 Instructon for Atomic Fact Chain Generation
As discussed before, we use the following prompt to decide whether one fact is entailed by others.
Instruction: Given a hypothesis and a reference, determine if the hypothesis is entailed by the reference. Entailed means that if we use the information from reference, we can infer the hypothesis, e.g. the hypothesis is in a logical relationship with the reference. If the hypothesis is entailed by the reference, answer True. If the hypothesis contradicts the reference or the hypothesis is neural,answer False. If the answer is ambiguous, answer False. The output should be True and False only. You should not explain why you made this choice. Now let’s begin the question:
Domonstrations
Example 1 Hypothesis: The cat is on the mat. Reference: The cat is sleeping on the mat. Answer: True Example 2 Hypothesis: The cat is on the mat. Reference: The dog and the rat are on the mat. Answer: False
Table 9: The instruction for atomic fact chain generation
A.3 Instruction for Document-level Relation Extraction
We use the following prompt with GPT-4 to conduct Document-level Relation Extraction.
Instruction: Given a text and entities in the given text, your task is to extract document-level relation triples from the text. A relation triple is a 3-tuple (subject, relation, object) where subject and object are named entities and relation is the relation between them. For example, given the text "Bill Gates is the founder of Microsoft", the extracted relation triple would be (Bill Gates, founder, Microsoft). In document-level RE, subject and object entities in a given triple might be dispersed across distinct sentences, and certain entities may have aliases in the form of distinct entity mentions. Therefore, you should pay attention to relations across sentences and entity aliases rather than sentence-level relations. i.e. You should not respond with sentence-level relations, but document-level relations. The answer format is listed below, you should obey it for any situation. Any additional information should not be added. Neither head entity nor tail entity should be empty strings. The relation triples should be in the following format: (’head entity’, ’relation’, ’tail entity’) (’head entity’, ’relation’, ’tail entity’) (’head entity’, ’relation’, ’tail entity’)
Demonstrations
Text
[…]
Entities
[…]
(’Bucharest’, ’country’, ’Romania’)
(’Bucharest’, ’located in the administrative territorial entity’, ’Romania’)
(’Romanian Air Force’, ’country’, ’Romania’)
(’Aurel Vlaicu Airport’, ’country’, ’Romania’)
(’Aurel Vlaicu Airport’, ’located in the administrative territorial entity’, ’Bucharest’)
(’Henri Coandă International Airport’, ’country’, ’Romania’)
(’Urziceni’, ’country’, ’Romania’)
(’Radu Berceanu’, ’country of citizenship’, ’Romania’)
[…]
(’Alexeni’, ’country’, ’Romania’)
Table 10: The instruction for Document-level Relation Extraction
A.4 Instruction for LLM validation
We use the prompt below to instruct LLMs focus on one piece of information each time.
Instruction: Given a hypothesis and a reference, determine if the hypothesis is entailed by the reference. Entailed means that if we use the information from reference, we can infer the hypothesis, e.g. the hypothesis is in a logical relationship with the reference. The hypothesis may be part of a logical chain, so we may provide you with the correctness of previous states. If we provide you with the correctness of previous states, you should not focus on the information that is redundant, that is, the information that exists both in previous facts and the current fact. Whether they are correct or not, you should not care them. You should only concentrate on new information. If the hypothesis is entailed by the reference, answer True. If the hypothesis contradicts the reference or the hypothesis is neural,answer False. If the answer is ambiguous, answer False. The output should be True and False only. You should not explain why you make this choice.
Table 11: The instruction for LLM validation
A.5 Instruction for LLM-based Evaluation of Informativenss
We use the following prompt in Table 12 to instruct LLMs to generate the informative score.
Instruction:
Given a Reference Answer and a Candidate Answer, your task is to first extract all information important to the Reference Answer and the Candidate Answer, and then calculate the Precision, Recall, and F1-score of the Candidate Answer Information pool concerning the Reference Answer Information pool. Information is defined as words, phrases or sentences that are important to the Reference Answer and the Candidate Answer. The subjective, objective, adverbial and attribute may all be included in the information. The information pool is the collection of all information extracted from the Reference Answer and the Candidate Answer. You can use the following function to calculate the Precision, Recall, and F1-score of the Candidate Answer Information pool with respect to the Reference Answer Information pool. def calculate_metrics(gold_key_information_pool, model_key_information_pool, matching_pool):
    precision = len(matching_pool) / len(gold_key_information_pool)
    recall = len(matching_pool) / len(model_key_information_pool)
    f1_score = 2 * precision * recall / (precision + recall)
    return f1_score
Finally, only return with a floating point number rounded to 3 decimal places.
Demonstrations
Reference Summary
background: coronary computed tomography angiography ( ccta ) is a frequently performed examination for coronary artery disease . when performed with retrospective gating , there is an opportunity to derive functional parameters of left ventricle utilizing automated software . complementary information , if validated with established standards , will enhance the total value of study.objective:study evaluates the usefulness of fully automated software for the assessment of left ventricular ejection fraction ( lvef ) using 64-slice ccta data and to correlate ct results with echocardiography ( echo ) . role of ct derived lv function is reviewed in the light of emerging technologies and recent developments in multidetector ct ( mdct).materials and methods : a total of 113 patients referred for mdct ccta for evaluation of coronary artery disease . all patients were scanned on 64 slice ge - helical ct scanner and had an echo done within 1 week of the ct scan . retrospectively electrocardiogram ( ecg)-correlated image reconstruction was performed with the reconstruction at 10% r - r interval increment . axial image sets were analyzed with advanced workstation using a program - auto ejection fraction , circulation : ge medical solutions.results:the mean lvef calculated by clinical echo was 58.6 4.5% and by fully automated software based on cta data was 58.9 5.4% . the pearson ’s regression analysis showed a large correlation , with a correlation coefficient of 0.503 ( p < 0.001 ) . bland - altman analysis showed a trend towards mdct resulting in slightly higher values for lvef when compared with echo.conclusion:the fully automated software is simple , reliable , and user - friendly , and can provide rapid assessment of lv functional parameters with good reproducibility . despite of good correlation , fewer patients are likely to benefit , in future , from this function due to smaller number of patients undergoing ccta with retrospective gating Candidate Summary
The text summarizes the results of a study on the use of multidetector computed tomography (MDCT) for assessing left ventricular (LV) function in patients undergoing coronary angiography. The study found that MDCT can provide useful functional information in addition to coronary angiography, and that fully automated analysis software can be used to measure LV ejection fraction (EF) with good reproducibility. However, the study also noted that the radiation dose involved in MDCT is a limitation, and that additional clinically indicated coronary CT examination with a lowest possible radiation dose is invaluable in patients known or suspected of ischemic heart disease. The study also found that validating consistency of results with magnetic resonance imaging (MRI) will further lend ‘ to the use of MDCT derived results. Thought Gold key information pool = [coronary computed tomography angiography, examination, coronary artery disease, retrospective gating,derive functional parameters,utilizing automated software, validated with established standards,enhance, total value, evaluates the usefulness, fully automated software, left ventricular ejection fraction, 64-slice ccta data, ct results, echocardiography ( echo ), reviewed, emerging technologies,multidetector ct ( mdct), 113 patients, evaluation, scanned,helical ct scanner, 1 week of,electrocardiogram, performed with the reconstruction, 10%r - r interval increment, axial image,analyzed with advanced workstation,auto ejection fraction,the mean, 58.6 4.5%, cta data was 58.9 5.4%, pearson ’s regression analysis,large correlation,coefficient of 0.503 ( p < 0.001 ), altman analysis,slightly higher values for lvef, compared, friendly,rapid assessment, good reproducibility, fewer patients, reliable, lvef] Model Key information pool = [multidetector computed tomography (MDCT), assessing left ventricular (LV), patients, coronary angiography, provide,useful functional information, automated analysis software, Measure, ejection fraction (EF), good reproducibility, radiation dose, limitation, clinically indicated, coronary CT examination, lowest possible, radiation dose, invaluable, patients known,suspected, ischemic heart disease, validating consistency, magnetic resonance imaging (MRI), further lend support, derived] Matching pool: [multidetector computed tomography (MDCT), assessing left ventricular (LV) match assessing left ventricular (LV), patients match patients, coronary angiography match coronary angiography, coronary angiography match coronary angiography, ejection fraction match lvef, good reproducibility match good reproducibility, slightly higher values match invaluable, derive functional parameters match derived]
F1 = calculate_metrics(gold_key_information_pool, model_key_information_pool, matching_pool)
print(F1) Output:
0.228
Table 12: The instruction for Document-level Relation Extraction
Appendix B Case Study
To provide a more intuitive and straightforward insight into the pipeline of EVA-Score, we provide a detailed case. From the case, we can see that EVA-Score provides not only a quantitative and objective information score but also provides a human-like explanation.
Reference Summary:
The mpp decreased the number of dopaminergic neurons by 40% , and increased the release of lactate dehydrogenase ( ldh ) into the culture medium . The tq significantly rescued dopaminergic neurons and decreased the release of ldh at the concentrations of 0.1 and 1 m
Atomic Fact Chains
[- The mpp decreased dopaminergic neurons, - The mpp decreased the number of dopaminergic neurons, - The mpp decreased the number of dopaminergic neurons by 40%]
[- The mpp increased the release of lactate dehydrogenase ( ldh ), - The mpp increased the release of lactate dehydrogenase ( ldh ) into the culture medium]
[- The tq significantly rescued dopaminergic neurons]
[- The tq decreased the release of ldh, - The tq decreased the release of ldh at the concentrations, - The tq decreased the release of ldh at the concentrations of 0.1 and 1 m]
Document-level Relations:
tq causes Thymoquinone
mpp leads to 1-methyl-4-phenylpyridinium
Candidate Summary:
The MPP significantly decreased the survival of dopaminergic neurons and increased the release of lysosomal hydrolase (LDH) into the culture medium. The neurotoxicity of MPP was attributed to its selective uptake by dopaminergic neurons through the dopamine transporter.
Atomic Fact Chains
[- The MPP significantly decreased dopaminergic neurons, -The MPP significantly decreased the survival of dopaminergic neurons ]
[- The MPP increased the release of lysosomal hydrolase (LDH), - The MPP increased the release of lysosomal hydrolase (LDH) into the culture medium]
[- The neurotoxicity was attributed to dopaminergic neurons, - The neurotoxicity of MPP was attributed to dopaminergic neurons, - The neurotoxicity of MPP was attributed to its selective uptake by dopaminergic neurons, - The neurotoxicity of MPP was attributed to dopaminergic neurons through the dopamine transporter, - The neurotoxicity of MPP was attributed to its selective uptake by dopaminergic neurons through the dopamine transporter.]
Document-level Relations:
mpp causes 1-methyl-4-phenylpyridinium
neurotoxicity infects neuroprotective effect
Matching Process
- “The mpp decreased dopaminergic neurons” match “The MPP significantly decreased dopaminergic neurons”
- “The mpp increased the release of lactate dehydrogenase ( ldh )” match “The mpp increased the release of lactate dehydrogenase ( ldh )”
- “The mpp increased the release of lactate dehydrogenase ( ldh ) into the culture medium” match “The MPP increased the release of lysosomal hydrolase (LDH) into the culture medium”
- “The mpp decreased dopaminergic neurons and increased the release of lactate dehydrogenase ( ldh )” match ‘The MPP significantly decreased dopaminergic neurons and increased the release of lysosomal hydrolase (LDH)”
- “The mpp decreased dopaminergic neurons and increased the release of lactate dehydrogenase ( ldh ) into the culture medium” match “The MPP significantly decreased dopaminergic neurons and increased the release of lysosomal hydrolase (LDH) into the culture medium”
- “mpp causes 1-methyl-4-phenylpyridinium” match “mpp leads to 1-methyl-4-phenylpyridinium”
Table 13: Case Study For EVA-Score from Pubmed Dataset.